# Reinforcing the Diffusion Chain of Lateral Thought with Diffusion Language Models

### v1

@@ Appendix E Ablation on Various Block Lengths
Table 8: Model accuracies with different block length on GSM8K dataset. LLaDA + DCoLT is trained with a generation length of 256 and 64 steps of reverse diffusion under respective block lengths.
@@
LLaDA + DCoLT 8B 83.17 83.11 82.27
LLaDA [27] divides the sequence into several blocks and generates responses blockwise from left to right in a semi-autoregressive manner. This policy introduces additional priors on the block-wise order of text generation. In Table. 8, for the baseline LLaDA 8B model, setting block length to $8$ achieves an accuracy of 63.26% on GSM8K, significantly higher than 49.47% without dividing blocks.
However, when we apply DCoLT to LLaDA with different block lengths, all models achieve similar performance, with the one with a block length 8 achieving slightly higher accurcy (83.17%) than the one with block length 256 without dividing blocks (82.27%). These results suggest that DCoLT does not rely on the blocking prior as in the baseline model.
@@ Appendix F Licenses for existing assets
The code and models associated with this paper will be released to the public later. For all code and data assets, we refer to their licenses in Table 9.

### v2

@@ Appendix E Ablation on Various Block Lengths
Table 8: Model accuracies with different block length on GSM8K dataset. LLaDOU (LLaDA + DCoLT) is trained with a generation length of 256 and 64 steps of reverse diffusion under respective block lengths.
@@
LLaDOU (LLaDA+DCoLT) 8B 83.17 83.11 82.27
LLaDA [27] divides the sequence into several blocks and generates responses blockwise from left to right in a semi-autoregressive manner. This policy introduces additional priors on the block-wise order of text generation. In Tab. 8, for the baseline LLaDA 8B model, setting block length to $8$ achieves an accuracy of 63.26% on GSM8K, significantly higher than 49.47% without dividing blocks.
However, when we apply DCoLT to LLaDA with different block lengths, all LLaDOU models achieve similar performance, with the one with a block length 8 achieving slightly higher accurcy (83.17%) than the one with block length 256 without dividing blocks (82.27%). These results suggest that LLaDOU does not rely on the blocking prior as in the baseline model.
@@ Appendix F Licenses for existing assets
The code and models associated with this paper will be released to the public later. For all code and data assets, we refer to their licenses in Tab. 9.